Solve for the equation of the family of curves in which the slope is (5βˆ’π‘₯)

(π‘¦βˆ’3)


Determine the type of the curve and the equation of a member passing through

(2 βˆ’ 1).



1
Expert's answer
2021-06-23T17:35:54-0400

"\displaystyle
\frac{\mathrm{d}y}{\mathrm{d}x} = \frac{5 - x}{y - 3} \\

y - 3 \mathrm{d}y = 5 - x \mathrm{d}x\\

\int y - 3 \mathrm{d}y = \int 5 - x \mathrm{d}x\\

\frac{y^2}{2} - 3y = 5x - \frac{x^2}{2} + C\\

\textsf{At}\,\, (2, -1)\\

\frac{(-1)^2}{2} - 3(-1) = 5(2) - \frac{2^2}{2} + C\\

\frac{7}{2} = 8 + C\\

C = -\frac{9}{2}\\

\frac{y^2}{2} - 3y = 5x - \frac{x^2}{2} -\frac{9}{2}\\

x^2 + y^2 - 10x - 6y + 9 = 0\\

(x - 5)^2 + (y - 3)^2 = 25 + 9 - 9 = 25\\

\textsf{This type of curve passing}\\
\textsf{through}\,\,(2,-1)\,\, \textsf{is a circle}\\\textsf{whose centre is}\,\, (5,3)\\
\textsf{and radius is}\,\,5."